solve the equation 4[p-2-(4+2p)] = -1 for P
Answer by jim_thompson5910(35256)

4[p-2-(4+2p)] = -1

4[p-2-4-2p] = -1

4[-p-6] = -1

-4p-24 = -1

-4p = -1+24

-4p = 23

p = 23/(-4)

p = -23/4